unl 100 60 Ho Hoo 200 (units of the ood

A $30 per unit production subsidy would cost how much to taxpayers?

$30,000
$24,000
$27,000

$26,500

A $30 per unit production tax would generate how much tax revenue?

$30,000
$24,000
$27,000
$26,500


A $30 per unit subsidy given to the producer would result in what new market price for consumers?

$20
$10
$40
$30

A $30 per unit tax on the producers would result in what new market price for consumers?

$60
$50
$70
$55

How much consumer surplus is shown?

None of these is correct
$24,000
$18,000
$16,000
$32,000

How much producer surplus is shown?

$18,000
$24,000
$32,000
None of these is correct
$16,000

What is the lowest level of per unit tax on producers that would result in the "disappearance" of this market (meaning that the price is so high that quantity demanded is driven to 0)?

$80
None of these answers is correct
$200
$60
$120
Answer #1

unl 60 s +subsid>y Ho Hoo (units of the ood.

2. Tax revenue = (40 - 10) x 800 = 30 x 800 = $ 24,000

3. Market price = $ 20

Because new supply curve and initial demand curve meets at price of $ 20.

4. New market price after tax = $ 60 because supply curve shift leftwards.

5. Consumer surplus is the area below demand curve and above market price.

C.S = 1/2 x 800 x (120 - 40) = 400 x 80 = $ 32000

6. Producer surplus is the area below market price and above supply curve.

P.S = 1/2 x 800 x (40 - 0) = 400 x 40 = $ 16000

7. Tax of $ 120 leads to quantity demanded to zero.